# Warped products and synthetic lower curvature bounds: an overview

### v1

@@ Abstract.
This is a survey about the contruction of warped products between (semi-)Riemannian manifolds and metric (measure) spaces. The resulting spaces will be semi-Riemannian manifolds, metric (measure) spaces or Lorentzian metric and metric measure spaces. We present details of the contruction in each case and we will highlight important properties like fiber independence and the energy equation. Warped products behave nicely in relation with curvature lower bounds. Here we will focus on sectional and Ricci curvature lower bounds and their Lorentzian counterparts. Throughout the article we provide many examples and formulate questions and conjectures.
@@
Warped products are a natural generalization of products of pairs of spaces. For instance, given two Riemannian manifolds $B$ and $F$, the base and the fiber, a warped product of $B$ and $F$ is obtained by the product space $B\times F$ equipped with a Riemannian metric that arises from the product metric by scaling the Riemannian metric $g_{\scriptscriptstyle{F}}$ on each fiber $\{r\}\times F$ according to a factor given by $f(r)>0$ for a smooth function $f$ on $B$. Warped products play an important role in Riemannian geometry. They give a wide class of examples of spaces with lower (and upper curvature bounds), including sectional, Ricci and Scalar curvature, and they arise as special spaces in rigidity theorems under lower curvature bounds. A warped product is a particular case of a Riemannian submersion and they can serve as building blocks in gluing constructions.
The construction also works for semi-Riemannian manifolds and in the case of a $1$-dimensional base $B$ one obtains a generalized cone that is a Lorentzian manifold. Prominent examples of warped products that satisfy the Einstein equation are the Robertson-Friedman cosmological model or the Schwartzschild solution that models the gravitational field of a star.
Finally warped products can be constructed in the context of metric and metric measure spaces. This yields new examples of non-smooth metric and metric measure spaces as well as Lorentzian length spaces.
In this paper we will survey the construction of warped products in smooth and non-smooth context, present different examples, explain crucial properties of warped products and also highlight the connection to synthetic curvature bounds.
@@ 2.3.4. Alexandrov differential
The next theorem generalizes the statement on sectional curvature bounds for warped products in the case of index 00 to the context of Alexandrov spaces with curvature bounded from below.
@@ 3.1. Warped products as Lorentzian length spaces
where we take the infimum w.r.t. all partitions $t_{0}=a\leq t_{1}\leq\dots\leq t_{N}=b$ of $[a,b]$. We note that by Lipschitz continuity of $f$, we have that $f\circ\gamma$ is absolutely continuous. Therefore this is infimum is the same as in [AGKS23, Definition 3.9].
Since $\gamma$ is causal one can show (see [AGKS23]) that
$|\alpha(t_{i})-\alpha(t_{i-1})|^{2}-(f\circ\alpha(t_{i-1}))^{2}\operatorname{d}_{\scriptscriptstyle{F}}(\beta(t_{i-1}),\beta(t_{i}))\geq 0.$

### v2

@@ Abstract.
This is a survey about the construction of warped products between (semi-)Riemannian manifolds and metric (measure) spaces. The resulting spaces will be semi-Riemannian manifolds, metric (measure) spaces or Lorentzian metric and metric measure spaces. We present details of the contruction in each case and we will highlight important properties like fiber independence and the energy equation. Warped products behave nicely in relation with curvature lower bounds. Here we will focus on sectional and Ricci curvature lower bounds and their Lorentzian counterparts. Throughout the article we provide many examples and formulate questions and conjectures.
@@
Warped products are a natural generalization of products of pairs of spaces. For instance, given two Riemannian manifolds $B$ and $F$, the base and the fiber, a warped product of $B$ and $F$ is obtained by the product space $B\times F$ equipped with a Riemannian metric that arises from the product metric by scaling the Riemannian metric $g_{\scriptscriptstyle{F}}$ on each fiber $\{r\}\times F$ according to a factor given by $f(r)>0$ for a smooth function $f$ on $B$.
Warped products play an important role in Riemannian geometry. They give a wide class of examples of spaces with lower and upper curvature bounds, including sectional, Ricci and Scalar curvature. They arise as special spaces in rigidity and almost rigidity theorems under lower curvature bounds. And they are a particular case of a Riemannian submersion and can serve as building blocks in gluing constructions, e.g. [Wra98].
Euclidean cones and suspensions are special cases of warped products of metric spaces that were known to respect curvature bounds in the sense of Alexandrov [BBI01, BGP92]. Famously Cheeger and Colding have shown that Riemannian manifolds satisfying certain lower Ricci curvature bounds and almost extremal volume or diameter conditions are Gromov–Hausdorff close to warped products of intervals with intrinsic metric spaces [CC96, CC97, Col96]. Chen and Alexander-Bishop then studied general warped products of intrinsic metric spaces and their relation to Alexandrov curvature bounds [Che99, AB98, AB04, AB05].
The warped product construction also works for semi-Riemannian manifolds and in the case of a $1$-dimensional base one obtains a generalized cone that is a Lorentzian manifold. Prominent examples of warped products that satisfy the Einstein equation are the Friedmann-Lemaître-Robertson-Walker cosmological models or the Schwarzschild solution that models the gravitational field of a star [O’N83, Sch16]. Finally the construction of warped products between metric and metric measure spaces yields new examples of non-smooth Lorentzian length spaces in the sense of Kunzinger and Sämann [KS18].
In this paper we will survey the construction of warped products in smooth and non-smooth context, present examples, explain crucial properties of warped products and highlight the connection to synthetic sectional and Ricci curvature lower bounds. Recently warped products also appeared in the context of lower Scalar curvature bounds (for instance [Gro18, CZ24]) but this will not be addressed in this article.
@@ 2.3.4. Alexandrov differential
The next theorem generalizes the statement on sectional curvature bounds for warped products in the case of signature (+,,+)(+$\dots$,+)( + , … , + ) to the context of Alexandrov spaces with curvature bounded from below.
@@ 3.1. Warped products as Lorentzian length spaces
where we take the infimum w.r.t. all partitions $t_{0}=a\leq t_{1}\leq\dots\leq t_{N}=b$ of $[a,b]$.